# KROMA: Ontology Matching with Knowledge Retrieval and Large Language Models

### v1

@@ Abstract
Ontology Matching (OM) is a cornerstone task of semantic interoperability, yet existing systems often rely on handcrafted rules or specialized models with limited adaptability. We present KROMA, a novel OM framework that harnesses Large Language Models (LLMs) within a Retrieval-Augmented Generation (RAG) pipeline, to dynamically enrich the semantic context of OM tasks with structural, lexical, and definitional knowledge. To optimize both performance and efficiency, KROMA integrates a bisimilarity-based concept matching and a lightweight ontology refinement step, which prune candidate concepts and substantially reduce the communication overhead from invoking LLMs. Through experiments on multiple benchmark datasets, we show that integrating knowledge retrieval with context-augmented LLMs significantly enhances ontology matching—outperforming both classic OM systems and cutting-edge LLM-based approaches—while keeping communication overhead comparable. Our study highlights the feasibility and benefit of the proposed optimization techniques (targeted knowledge retrieval, prompt enrichment, and ontology refinement) for ontology matching at scale. Code is available at: https://anonymous.4open.science/r/kroma/
@@ 1 Introduction
Ontologies have been routinely developed to unify and standardize knowledge representation to support data-driven applications. They allow researchers to harmonize terminologies and enhance knowledge and sharing in their fields. Ontologies can be classified according to their level of specificity, ranging from more abstract, general-purpose ontologies such as Basic Formal Ontology (BFO)[38] or DOLCE[6] down to more domain-oriented ‘mid-level” ones, such as CheBi [14] in chemistry, Industrial Ontology Foundy (IOF) [17] or Common Core Ontology(CCO) [27]. Domain ontologies are data-driven, task-specific “low-level” ontologies, containing concepts from domain-specific data, such as Materials Data Science Ontology (MDS-Onto) [46]. To achieve broad usability, ontologies need to be aligned for better interoperability via ontology matching.
Ontology matching has been studied to find correspondence between terms that are semantically equivalent [52]. It is a cornerstone task to ensure semantic interoperability among terms originated from different sources. Ontology matching methods can be categorized to rule-based or structural-based (graph pattern or path-based) matching [10], matching with semantic similarity, machine learning-based approaches and hybrid methods. Linguistic (terminological) methods are often used for ontology matching tasks, ranging from simple string matching or embedding learning to advance counterparts based on natural language processing (NLP). Nevertheless, conventional rule- or structural-based methods are often restricted to certain use cases and hard to be generalized for new or unseen concepts. Learning-based approaches may on the other hand require high re-training process, for which abundant annotated or training data remains a luxury especially for e.g., data-driven scientific research.
@@
This paper introduces $\mathsf{KROMA}$, a novel framework that exploits $\mathsf{LLMs}$ to enhance ontology matching. Unlike existing $\mathsf{LLM}$-based methods that typically “outsource” ontology matching to $\mathsf{LLMs}$ with direct prompting (which may have low confidence and risk of hallucination), $\mathsf{KROMA}$ maintain a set of conceptually similar groups that are co-determined by concept similarity and $\mathsf{LLMs}$, both guided by their “context” obtained via a runtime knowledge retrieval process. Moreover, the groups are further refined by a global ontological equivalence relation that incorporate structural equivalence.
Contributions. Our main contributions are summarized below.
@@
(2) We introduce $\mathsf{KROMA}$, an $\mathsf{LLM}$-enhanced ontology matching framework (Section 4). $\mathsf{KROMA}$ fine-tunes $\mathsf{LLMs}$ with prompts over enriched semantic contexts. Such contexts are obtained from knowledge retrieval, referencing high-quality, external knowledge resources.
(3) $\mathsf{KROMA}$ supports both offline and online matching, to “cold-start” from scratch, and to digest terms arriving from a stream of data, respectively. We introduce efficient algorithms to correctly construct and maintain the optimal concept graphs (Section 5). (a) The offline refinement performs a fast grouping process guided by the bisimilar equivalence, blending concept equivalence tests co-determined by semantic closeness and $\mathsf{LLMs}$. (b) The online algorithm effectively incrementalizes its offline counterpart with fast delay time for continuous concept streams. Both algorithms are in low polynomial time, with optimality guarantee on the computed concept graphs.
(4) Using benchmarking ontologies and knowledge graphs, we experimentally verify the effectiveness and efficiency of $\mathsf{KROMA}$ (Section 6). We found that $\mathsf{KROMA}$ outperforms existing $\mathsf{LLM}$-based methods by 10.95% on average, and the optimization of knowledge retrieval and refinement improves its accuracy by 6.65% and 2.68%, respectively.
Related Work. We summarize related work below.
@@ 2 Ontologies and Ontology Matching
$[c]\in V_{O}$ an equivalence class in $\mathcal{C}$
@@
Ontologies. An ontology $O$ is a pair $(C,E)$, where $C$ is a finite set of concept names, and $E\subset C\times C$ is a set of relations. An ontology has a graph representation with a set of concept nodes $C$, and a set of edges $E$. We consider ontologies as directed acyclic graphs (DAGs).
In addition, each concept node (or simply “node”) $c$ in $O$ carries the following auxiliary structure. (1) The rank of a node $c\in C$ is defined as: (a) $r(c)$ = 0 if $c$ has no child, otherwise, (b) $r(c)$ = $\max(r(c^{\prime})+1)$ for any child $c^{\prime}$ of $c$ in $O$. (2) A ground set $c.{\mathcal{I}}$, refers to a set of entities from e.g., external ontologies or knowledge bases that can be validated to belong to the concept $c$.
Ontology Matching. Given a source ontology $O_{s}$ = $(C_{s},E_{s})$ and a target ontology $O_{t}$ = $(C_{t},E_{t})$, an ontological equivalence relation $R_{\simeq}\subseteq C_{s}\times C_{t}$ is an equivalence relation that contains pairs of nodes $(c,c^{\prime})$ that are considered to be “semantically equivalent”. The relation $R_{\simeq}$ induces an equivalence partition ${\mathcal{C}}$ of the concept set $C_{s}\cup C_{t}$, such that each partition is an equivalence class that contains a set of pairwise equivalent concepts in $C_{s}\cup C_{t}$.
Consistently, we define a concept graph ${\mathcal{G}}_{O}$ = $(V_{O},E_{O})$ as a DAG with a set of nodes $V_{O}$, where each node $[c]\in V_{O}$ is an equivalence class in ${\mathcal{C}}$, and there exists an edge $([c],[c^{\prime}])\in E_{O}$ if and only if there exists an edge $(c,c^{\prime})$ in $E_{s}$ or in $E_{t}$. A concept graph ${\mathcal{G}}_{O}$ can be a multigraph: there may exist multiple edges of different relation names between two equivalent classes.
Given $O_{s}$ and $O_{t}$, the task of ontology matching is to properly formulate $R_{\simeq}$ and compute ${\mathcal{C}}$ induced by $R_{\simeq}$ over $O_{s}$ and $O_{t}$; or equivalently, compute the concept graph ${\mathcal{G}}_{O}$ induced by $R_{\simeq}$.
@@
Figure 1 depicts two ontologies $O_{s}$ and $O_{t}$ as DAGs, involving in total 9 concept nodes. An equivalence relation may suggest that “mammal”, “animal”, “organism” and “vertebrate” are pairwise similar; and similarly for the sets {“house pet”, “carnivora” and “canine”}, and {“wolfdog, “coyote”}”. This induces a concept graph ${\mathcal{G}}_{o}$ as shown on the top right as a result of ontology matching, with three equivalence classes.
@@
We specify a prompt query $q$ and $\mathsf{LLMs}$ for ontology matching. A prompt query $q$ is in the form of $q(c,c^{\prime})$, which asks “are $c$ and $c^{\prime}$ semantically equivalent?” An $\mathsf{LLM}$ $M$ can be queried by $q(c,c^{\prime})$ and acts as a Boolean “oracle” with “yes/no” answer. An $\mathsf{LLM}$ is deterministic, if it always generate a same answer for the same prompt query. We consider deterministic $\mathsf{LLMs}$, as in practice, such $\mathsf{LLMs}$ are desired for consistent and robust performance.
@@ 3.1 Semantic Equivalence: A Characterization
[…]
@@
where $q$ is a prompt query that specifies the context of concept equivalence for $\mathsf{LLMs}$, and $\gamma$ be a configurable parameter. $\mathsf{KROMA}$ supports a built-in library of semantic similarity functions $\mathsf{sim}$, including (a) string similarity, feature and information measure [40], or normalized distances (NGDs) [28]; and (b) a variety of $\mathsf{LLMs}$ such as GPT-4o Mini (OpenAI) [36], LLaMA-3.3 (Meta AI) [31], and Qwen-2.5 (Qwen Team) [43].
Ontological Bisimilarity. We next provide a specification of the ontological equivalence relation, notably, ontological bisimilarity, denoted by the same symbol $R_{\simeq}$ for simplicity. Given a source ontology $O_{s}$ = $(C_{s},E_{s})$, and a target ontology $O_{t}$ = $(C_{t},E_{t})$, we say a pair of nodes $c_{s}\in C_{s}$ and $c_{t}\in C_{t}$ are ontologically bisimilar, denoted as $(c_{s},c_{t})\in R_{\simeq}$, if and only if there exists a non-empty binary relation $R_{\simeq}$, such that: (1) $c_{s}$ and $c_{t}$ are concept similar, i.e., $F(c_{s},c_{t})\geq\alpha$, for a threshold $\alpha$; (2) for every edge $(c_{s}^{\prime},c_{s})\in E_{s}$, there exists an edge $(c_{t}^{\prime},c_{t})\in E_{t}$, such that $(c_{s}^{\prime},c_{t}^{\prime})\in R_{\simeq}$, and vice versa; and (3) for every edge $(c_{s},c_{s}^{\prime\prime})\in E_{s}$, there exists an edge $(c_{t},c_{t}^{\prime\prime})\in E_{t}$, such that $(c_{s}^{\prime\prime},c_{t}^{\prime\prime})\in R_{\simeq}$.
One can verify the following result.
@@
We can prove the above results by verifying that $R_{\simeq}$ is reflexive, symmetric and transitive over the concept set $C_{s}\cup C_{t}$, ensured by the transitivity of concept similarity and by definition. Observe that two nodes that are concept similar may not be ontologically bisimilar. On the other hand, two ontologically bisimilar entities must be concept similar, by definition.
@@ 3.2 Problem Statement
[…]
We justify the above characterization by proving that there exists an “optimal”, invariant solution encoded by a concept graph ${\mathcal{G}}_{O}$. To see this, we provide a minimality and uniqueness guarantee on ${\mathcal{G}}_{O}$.
@@
Given a configuration $W$ and semantic equivalence specified by the ontological bisimilar relation $R_{\simeq}$, there is a unique smallest concept graph ${\mathcal{G}}_{O}$ that capture all semantically equivalent nodes in terms of $R_{\simeq}$.
@@
The above analysis suggests that for a configuration $W$, it is desirable to compute such an optimal concept graph as an invariant, stable result with guarantees on sizes and uniqueness on topological structures. We next introduce $\mathsf{KROMA}$ and efficient algorithms to compute such optimal concept graphs.
@@ 4.1 Framework Overview
[…]
Concept Graph Initialization Upon receiving two ontologies $O_{s}$ = $(C_{s},E_{s})$ and $O_{t}$ = $(C_{t},E_{t})$, $\mathsf{KROMA}$ initializes the concept graph $\mathcal{G}_{O}$ = $(V,E)$ with $V=C_{S}\cup C_{T}$ and $E=E_{S}\cup E_{T}$. For each concept $c\in V$, $\mathsf{KROMA}$ computes the $rank$ $r(c)$ as described in Section 2.
Knowledge Retrieval. To assemble a rich, yet compact, context for each concept $c\in V$, we perform: (1) Neighborhood Sampling: traverse up to its two hops in $\mathcal{G}_{O}$ to collect parents, children, and “sibling” concepts of $c$, creating a node induced subgraph centered at $c$; (2) subgraph parameterization: Sample and substitute constant values from the subgraph with variables to create SPARQL queries $S_{q}$; and (3) ground set curation: process $S_{q}$ onto external knowledge bases to augment its ground set $c.{\mathcal{I}}$ with auxiliary information (e.g., relevant entities, definition, labels, etc.).
Embedding Generation. In this phase, $\mathsf{KROMA}$ consults a built-in library of semantic similarity functions ${\mathsf{sim}}(\cdot,\cdot)$ and embedding functions ${\mathsf{embd}}(\cdot)$. For each concept $c\in V$, $\mathsf{KROMA}$ computes:
@@
where ${\mathsf{embd}}_{graph}$ (e.g. node2vec [22]) captures the $c$’s topology information and ${\mathsf{embd}}_{text}$ (e.g. SciBERT [5]) encodes $c$’s textual context. After obtaining the necessary embeddings ($z_{c}$ $\forall c\in V$), $\mathsf{KROMA}$ computes pairwise concept similarity between source and target ontologies using $\mathsf{sim}$ functions:
@@
From $\mathbb{S}$, we select the top-$k$ pairs with the highest similarity scores (i.e. in descending order of $score_{s,t}$), yielding the candidate list $\mathbb{C}$.
@@
We revisit Example 1. (1) A knowledge retrieval for node “house pet” samples its neighbors and issues a set of “star-shaped” SPARQL queries centered at “house pet” to query an underlying knowledge graph $KG$. This enriches its ground set with a majority of herbivore or omnivorous pets that are not “carnivore”. Similarly, the ground set of “carnivore” is enriched by “wolf”, unlikely a house pet. Despite “coyote” and “wolfdog” (house pet) alone are less similar, the ground set of “coyote” turns out to be a set of canine pets e.g., “husky” that are “coyote-like”, hence similar with “wolfdog”. (2) The embedding generation phase incorporates enriched ground sets and generate embeddings accordingly, which scores that distinguishes “house pet” from “carnivore” due to embedding difference, and assert “coyote” and “wolfdog” to be concept similar, hence a candidate pair to be “double checked” by $\mathsf{LLMs}$ in the next phase.
[…]
@@
Ontology Refinement & Expert Validation. $\mathsf{KROMA}$ next invokes a refinement process, $\mathsf{OfflineRefine}$ (Algorithm 1), to “cold-start” the construction of ${\mathcal{G}}_{O}$, or $\mathsf{OnlineRefine}$ (Algorithm 2), to incrementally refine $\mathcal{G}_{O}$ for any unseen, newly arrived concept nodes or edges. Any pair of nodes whose structural ranks remain in “conflict” is routed into a set of queries for expert validation (see Algorithm 2, Section 4); once approved, are integrated back into ${\mathcal{G}}_{O}$. An active sampling strategy is applied, to select pairs of nodes that have low confidence from $\mathsf{LLMs}$ for expert validation, to reduce the manual effort.

### v2

@@ Abstract
Ontology Matching (OM) is a cornerstone task of semantic interoperability, yet existing systems often rely on handcrafted rules or specialized models with limited adaptability. We present KROMA, a novel OM framework that harnesses Large Language Models (LLMs) within a Retrieval-Augmented Generation (RAG) pipeline, to dynamically enrich the semantic context of OM tasks with structural, lexical, and definitional knowledge. To optimize both performance and efficiency, KROMA integrates a bisimilarity-based concept matching and a lightweight ontology refinement step, which prune candidate concepts and substantially reduce the communication overhead from invoking LLMs. Through experiments on multiple benchmark datasets, we show that integrating knowledge retrieval with context-augmented LLMs significantly enhances ontology matching—outperforming both classic OM systems and cutting-edge LLM-based approaches—while keeping communication overhead comparable. Our study highlights the feasibility and benefit of the proposed optimization techniques (targeted knowledge retrieval, prompt enrichment, and ontology refinement) for ontology matching at scale. Our code and experimental dataset has been made available at: https://github.com/lamng3/kroma
@@ 1 Introduction
Ontologies have been routinely developed to unify and standardize knowledge representation to support data-driven applications. They allow researchers to harmonize terminologies and enhance knowledge and sharing in their fields. Ontologies can be classified according to their level of specificity, ranging from more abstract, general-purpose ontologies such as Basic Formal Ontology (BFO) [35] or DOLCE [5] down to more domain-oriented ‘mid-level” ones, such as CheBi [14] in chemistry, Industrial Ontology Foundy (IOF) [17] or Common Core Ontology (CCO) [26]. Domain ontologies are data-driven, task-specific “low-level” ontologies, containing concepts from domain-specific data, such as Materials Data Science Ontology (MDS-Onto) [42]. To achieve broad usability, ontologies need to be effectively aligned for better interoperability using ontology matching.
Ontology matching has been studied to find correspondence between terms that are semantically equivalent [50]. It is a cornerstone task to ensure semantic interoperability among terms originated from different sources. Ontology matching methods can be categorized to rule-based or structural-based (graph pattern or path-based) matching [10], matching with semantic similarity, machine learning-based approaches and hybrid methods. Linguistic (terminological) methods are often used for ontology matching tasks, ranging from simple string matching or embedding learning to advance counterparts based on natural language processing (NLP). Nevertheless, conventional rule- or structural-based methods are often restricted to certain use cases and hard to be generalized for new or unseen concepts. Learning-based approaches may on the other hand require expensive re-training process, for which abundant annotated or training data remains a luxury especially for e.g., data-driven scientific research.
@@
This paper introduces $\mathsf{KROMA}$, a novel framework that exploits $\mathsf{LLMs}$ to enhance ontology matching. Unlike existing $\mathsf{LLM}$-based methods that typically “outsource” ontology matching to $\mathsf{LLMs}$ using direct prompting (which may result in low confidence and risk of hallucination), $\mathsf{KROMA}$ maintains a set of conceptually similar groups that are co-determined by concept similarity and language models, both guided by their “augmented context” obtained via a runtime knowledge retrieval process. Moreover, the groups are further refined by a global ontological equivalence relation that incorporate structural equivalence.
Contributions. Our main technical contributions are summarized below.
@@
(2) We introduce $\mathsf{KROMA}$, an ontology matching framework leveraging the power of $\mathsf{LLMs}$ (Section 4). $\mathsf{KROMA}$ fine-tunes $\mathsf{LLMs}$ with prompts over enriched semantic contexts. Such contextual information are obtained from the knowledge retrieval process, referencing high-quality, external knowledge resources.
(3) $\mathsf{KROMA}$ supports both offline and online matching, to “cold-start” from scratch, and to digest terms arriving from a stream of data, respectively. We introduce efficient algorithms to correctly construct and maintain the optimal concept graphs (Section 5). (a) Offline refinement algorithm performs a fast grouping process guided by the bisimilar equivalence property, blending the concept equivalence tests co-determined by semantic closeness and language models. (b) Online refinement algorithm effectively incrementalizes its offline counterpart with fast delay time for continuous concept streams. Both algorithms are in low polynomial time, with optimality guarantee on the computed concept graphs.
(4) Using benchmarking ontologies and knowledge graphs, we experimentally verify the effectiveness and efficiency of $\mathsf{KROMA}$ (Section 6). We found that $\mathsf{KROMA}$ outperforms existing methods using large languge models by 10.95% on average, and the optimization of utilizing knowledge retrieval and refinement processes improves $\mathsf{KROMA}$’s accuracy by 6.65% and 2.68%, respectively.
Related Work. We summarize the related work below.
@@ 2 Ontologies and Ontology Matching
$c]\in V_{O}$ an equivalence class in $\mathcal{C}$
@@
Ontologies. An ontology $O$ is a pair $(C,E)$, where $C$ is a finite set of concept names, and $E\subset C\times C$ is a set of relations between the concepts. An ontology has a graph representation with a set of concept nodes $C$, and a set of edges $E$. In this paper, we consider ontologies as directed acyclic graphs (DAGs).
In addition, each concept node (or simply “node”) $c$ in $O$ carries the following auxiliary structure. (1) The rank of a node $c\in C$ is defined as: (a) $r(c)$ = 0 if $c$ has no child, otherwise, (b) $r(c)$ = $\max(r(c^{\prime})+1)$ for any child $c^{\prime}$ of $c$ in $O$. (2) A ground set $c.{\mathcal{I}}$, refers to a set of auxilary entities from e.g., external ontologies or knowledge bases, that can be validated to belong to the concept $c$.
Ontology Matching. Given a source ontology $O_{s}$ = $(C_{s},E_{s})$ and a target ontology $O_{t}$ = $(C_{t},E_{t})$, an ontological equivalence relation $R_{\simeq}\subseteq C_{s}\times C_{t}$ is defined as an equivalence relation that contains pairs of nodes $(c,c^{\prime})$ that are considered to be “semantically equivalent”. The relation $R_{\simeq}$ induces an equivalence partition $\mathcal{C}$} of the concept set $C_{s}\cup C_{t}$, such that each partition is an equivalence class that contains a set of pairwise equivalent concepts in $C_{s}\cup C_{t}$.
Consistently, we define a concept graph ${\mathcal{G}}_{O}$ = $(V_{O},E_{O})$ as a DAG with a finite set of nodes $V_{O}$, where each node $c]\in V_{O}$ is an equivalence class in $\mathcal{C}$}, and there exists an edge $([c],[c^{\prime}])\in E_{O}$ if and only if there exists an edge $(c,c^{\prime})$ in $E_{s}$ or in $E_{t}$. A concept graph ${\mathcal{G}}_{O}$ can be a multigraph: there may exist multiple edges of different relation names between two equivalent classes.
Given $O_{s}$ and $O_{t}$, the task of ontology matching is to formulate $R_{\simeq}$ and $\mathcal{C}$} induced by $R_{\simeq}$ and compute the concept graph ${\mathcal{G}}_{O}$ accordingly.
@@
Figure 1 depicts two ontologies $O_{s}$ and $O_{t}$ as DAGs with 9 concept nodes. An equivalence relation may suggest that “mammal”, “animal”, “organism” and “vertebrate” are pairwise similar; and similarly for the sets {“house pet”, “carnivora” and “canine”}, and {“wolfdog, “coyote”}”. This induces a concept graph ${\mathcal{G}}_{o}$ as a result of ontology matching, with three equivalence classes.
@@
We specify a prompt query $q$ and a large language model $M$ for ontology matching. A prompt query $q$ is in the form of $q(c,c^{\prime})$, which asks “are $c$ and $c^{\prime}$ semantically equivalent?” An $\mathsf{LLM}$ $M$ can be queried by $q(c,c^{\prime})$ and acts as a Boolean “oracle” with “yes/no” answer. An $\mathsf{LLM}$ is deterministic, if it always generate a same answer for the same prompt query. We consider deterministic $\mathsf{LLMs}$, as in practice, such $\mathsf{LLMs}$ are desired for consistent and robust performance.
@@ 3.1 Semantic Equivalence: A Characterization
Concept similarity. A variety of methods have been proposed to determine whether two concepts are semantically equivalent [7]. $\mathsf{KROMA}$ by default uses a Boolean function $F$ defined by a weighted combination of a semantic closeness metric $\mathsf{sim}$} [^1] We adopt similarity metrics that satisfy transitivity, i.e., if concept c is similar to c’, and c’ is similar to c”, then c is similar to c”. This is to ensure transitivity of ontology equivalence, and is practical for representative concept similarity measures. and the result from a set of $\mathsf{LLMs}$ $\mathcal{M}$} (see Section 4).
@@
where $q$ is a prompt query that specifies the context of concept equivalence for $\mathsf{LLMs}$, and $\gamma$ be a configurable parameter. $\mathsf{KROMA}$ supports a built-in library of semantic similarity functions $\mathsf{sim}$, including (a) string similarity, feature and information measure [37], or normalized distances (NGDs) [27]; and (b) a variety of $\mathsf{LLMs}$ such as GPT-4o Mini [34], LLaMA-3.3 [21], and Qwen-2.5 [39].
Ontological Bisimilarity. We next provide a specification of the ontological equivalence relation, notably, ontological bisimilarity, denoted by the same symbol $R_{\simeq}$ for simplicity. Given a source ontology $O_{s}$ = $(C_{s},E_{s})$, and a target ontology $O_{t}$ = $(C_{t},E_{t})$, we say a pair of nodes $c_{s}\in C_{s}$ and $c_{t}\in C_{t}$ are ontologically bisimilar, denoted as $(c_{s},c_{t})\in R_{\simeq}$, if and only if there exists a non-empty binary relation $R_{\simeq}$, such that: (1) $c_{s}$ and $c_{t}$ are conceptually similar, i.e., $F(c_{s},c_{t})\geq\alpha$, for a threshold $\alpha$; (2) for every edge $(c_{s}^{\prime},c_{s})\in E_{s}$, there exists an edge $(c_{t}^{\prime},c_{t})\in E_{t}$, such that $(c_{s}^{\prime},c_{t}^{\prime})\in R_{\simeq}$, and vice versa; and (3) for every edge $(c_{s},c_{s}^{\prime\prime})\in E_{s}$, there exists an edge $(c_{t},c_{t}^{\prime\prime})\in E_{t}$, such that $(c_{s}^{\prime\prime},c_{t}^{\prime\prime})\in R_{\simeq}$.
@@
We can prove the above results by verifying that $R_{\simeq}$ is reflexive, symmetric, and transitive over the concept set $C_{s}\cup C_{t}$, ensured by the transitivity of concept similarity and by definition. Observe that two entities that are conceptually similar may not be ontologically bisimilar. On the other hand, two ontologically bisimilar entities must be conceptually similar, following the definition.
@@ 3.2 Problem Statement
We now formulate our ontological matching problem. Given a configuration $W$ = $(O_{s},O_{t},F,{\mathcal{M}},\alpha)$ that specifies as input a source ontology $O_{s}$, a target ontology $O_{t}$, a Boolean function $F$ and threshold $\alpha\in(0,1]$ that determines concept similarity, and a set of $\mathsf{LLMs}$ $\mathcal{M}$}, the problem becomes computing a smallest concept graph ${\mathcal{G}}_{O}$ induced by the ontologically bisimilar equivalence $R_{\simeq}$.
We can justify the above characterization by proving that there exists an “optimal”, invariant solution encoded by a concept graph ${\mathcal{G}}_{O}$. To arrive at this, we provide a minimality and uniqueness guarantee on ${\mathcal{G}}_{O}$ as below.
@@
Given a configuration $W$ and semantic equivalence specified by the ontological bisimilar relation $R_{\simeq}$ property, there is a unique smallest concept graph ${\mathcal{G}}_{O}$ that captures all semantically equivalent nodes in terms of $R_{\simeq}$.
@@
The above analysis suggests that for a configuration $W$, it is desirable to compute such an optimal concept graph as an invariant, stable result with guarantees on sizes and uniqueness on topological structures. We next introduce $\mathsf{KROMA}$ and efficient algorithms to compute the aforementioned optimal concept graphs.
@@ 4.1 Framework Overview
Given a configuration $W$ = $(O_{s},O_{t},F,{\mathcal{M}})$ where $\mathcal{M}$} is a set of pre-trained large language models $\mathcal{M}$, $\mathsf{KROMA}$ has the following major functional modules that enables a robust and effective multi-session ontology matching process.
Concept Graph Initialization Upon receiving two ontologies $O_{s}$ = $(C_{s},E_{s})$ and $O_{t}$ = $(C_{t},E_{t})$ as the inputs, $\mathsf{KROMA}$ initializes the concept graph $\mathcal{G}_{O}$ = $(V,E)$ with $V=C_{S}\cup C_{T}$ and $E=E_{S}\cup E_{T}$. For each concept $c\in V$, $\mathsf{KROMA}$ then globally computes the structural rankrank $r(c)$ as described in Section 2.
Knowledge Retrieval. To assemble a rich, yet compact, context for each concept $c\in V$, we perform: (1) Neighborhood Sampling: traverse up to its two hops in $\mathcal{G}_{O}$ to collect parents, children, and “sibling” concepts of $c$, creating a node induced subgraph centered at $c$; (2) Subgraph Parameterization: Sample and substitute constant values from the subgraph with variables to create SPARQL queries $S_{q}$; and (3) Ground Set Curation: Apply the generated SPARQL queries $S_{q}$ onto external knowledge bases to augment the concept’s ground set $c.{\mathcal{I}}$ with auxiliary information (e.g., relevant entities, definition, labels, synonyms, etc.).
Embedding Generation. In this phase, $\mathsf{KROMA}$ leverages a built-in library of semantic similarity functions ${\mathsf{sim}}(\cdot,\cdot)$ and embedding functions ${\mathsf{embd}}(\cdot)$. For each concept $c\in V$, $\mathsf{KROMA}$ computes the joint textual and structural embedding:
@@
where ${\mathsf{embd}}_{graph}$ (e.g. node2vec [22]) captures the $c$’s topology information and ${\mathsf{embd}}_{text}$ (e.g. SciBERT [4]) encodes $c$’s textual context. After obtaining the necessary embeddings ($z_{c}$ $\forall c\in V$), $\mathsf{KROMA}$ then computes pairwise concept similarity between source and target ontologies using the $\mathsf{sim}$ functions:
@@
From $\mathbb{S}$, we select the top-$k$ pairs with the highest similarity scores (i.e. in descending order of $score_{s,t}$), yielding the best candidate list $\mathbb{C}$ to ask $\mathsf{LLMs}$.
@@
We revisit Example 1. (1) A knowledge retrieval for node “house pet” samples its neighbors and issues a set of “star-shaped” SPARQL queries centered at “house pet” to query an underlying knowledge graph $KG$. This enriches its ground set with a majority of herbivore or omnivorous pets that are not “carnivore”. Similarly, the ground set of “carnivore” is enriched by “wolf”, unlikely a house pet. Despite “coyote” and “wolfdog” (house pet) alone are less similar, the ground set of “coyote” turns out to be a set of canine pets e.g., “husky” that are “coyote-like”, hence similar with “wolfdog”. (2) The embedding generation phase incorporates enriched ground sets and generate embeddings accordingly, which scores that distinguishes “house pet” from “carnivore” due to embedding difference, and assert “coyote” and “wolfdog” to be conceptually similar, allowing the candidate pair to be “double checked” by $\mathsf{LLMs}$ in the following phase.
Prompt Querying LLMs. For each candidate pair $(c_{s},c_{t})\in\mathbb{C}$, $\mathsf{KROMA}$ generates an NL prompt that includes: (1) Task description $\mathcal{T}$ (e.g. “Given two ontology concepts and their metadata, decide if they are related or not.”), (2) In-context examples $\mathcal{D}$ containing both positive and negative matches, (3) Query context for $c_{s}$ and $c_{t}$ including their ground sets, (4) Output format and confidence (e.g., “Answer Yes or No, and provide a confidence score between 0 and 10.”). $\mathsf{KROMA}$ then calls (a set of) $\mathsf{LLMs}$ $\mathcal{M}$ to obtain a matching decision with confidence. Low-confidence or conflicting outputs are routed to validation module.
A query template and a generated example is illustrated in Figure 3.
@@
Ontology Refinement & Expert Validation. $\mathsf{KROMA}$ next invokes a refinement process, $\mathsf{OfflineRefine}$ (Algorithm 1), to “cold-start” the construction of ${\mathcal{G}}_{O}$, or $\mathsf{OnlineRefine}$ (Algorithm 2), to incrementally refine $\mathcal{G}_{O}$ for any unseen, newly arrived concept nodes or edges. Any pair of nodes whose structural ranks remain in “conflict” is routed into a set of queries for expert validation (see Algorithm 2, Section 4); once approved, are integrated back into ${\mathcal{G}}_{O}$. An active sampling strategy is applied, to select pairs of nodes that have low confidence from $\mathsf{LLMs}$ for expert validation, with an aim to minimalize the manual effort needed.